# Critical dynamics of Model H within the real-time fRG approach

### v1

@@ II Langevin equation for Model H
In Eq. 1 and 2 $\lambda_{0}$ and $\eta_{0}$ are the transport coefficient and shear viscosity, respectively. Their dynamical scaling behaviors are described by the exponents $x_{\lambda}$ and $x_{\eta}$ in the critical regime, and these exponents can also be seen as dynamic anomalous dimensions. The Hamiltonian has the usual form except the term $\bm{j}^{2}/2$ included. Thus the static critical properties of $\phi$ are the same as the Ginzburg-Landau model with Ising symmetry. When the system is in thermal equilibrium, the equilibrium distribution function is given by $e^{-\mathscr{H}}$. The coupling constant $g_{0}$ describes the interaction between the two dynamical modes in model H. The Gaussian noises $\theta$ and $\zeta$ derive the system to thermal equilibrium and their correlation functions are shown in Eq. 4.
@@
$D=\frac{\lambda}{\chi_{\phi}}=\frac{R\,k_{B}T}{\eta\xi}\,,$ (7)
with the static susceptibility $\chi_{\phi}$. Consequently, one arrives at the dynamic critical exponent $z$ for the order parameter
@@
where $\eta_{\phi}$ is the static anomalous dimension. The results of the exponents $x_{\lambda}$ and $x_{\eta}$ in the $\epsilon$-expansion up to the order of $O(\epsilon^{2})$ are obtained as [53]
@@
In the next section, we will compute the above exponents using the functional renormalization group approach.
@@ III Model H within the real-time fRG approach
$\Gamma_{k}[\Phi]$ $=i\int\mathrm{d}^{d}x\mathrm{d}t\,\,\phi_{q}\left(\frac{\partial}{\partial t}\phi_{c}+\lambda_{k}\bm{\nabla}^{2}\bm{\nabla}^{2}\phi_{c}\right)$
@@
$-g_{k}\bm{\nabla}\phi_{c}(-\bm{\nabla}^{2}\phi_{c})+2\eta_{k}\bm{\nabla}^{2}\bm{j}_{\bm{q},\beta}\Bigg{]}\,.$ (10)
Here, $(\phi_{c},\phi_{q})$ and $(\bm{j}_{\bm{c},\beta},\bm{j}_{\bm{q},\alpha})$ are the order parameter fields and momentum densities with the subscripts ‘c’ and ‘q’ denoting classical and quantum fields, respectively, see [43, 47] for more details. And subscripts $\alpha,\beta$ denote the component of vector $\bm{j}$, i.e., $\alpha,\beta=1,2,...,d$. Different from the analogue in Eq. 1 and 2, the transport and coupling constants $\lambda_{k},\eta_{k}$ and $g_{k}$ in Eq. 10 also receive corrections from interactions, and they are RG-scale dependent. Note that the order parameter belongs to the Ising universality class, thus presents the same static critical properties. The order parameter can be regarded as a linear combination of the chiral condensate and baryon density in QCD [30]. Moreover, one can see that the powers of gradient in front of $\phi_{c}$ are higher than those in Model A or Model C due to the conservation of the order parameter and momentum. The higher order scatterings among the $\phi$ fields are encoded in the effective potential as well as the mass term of order parameter.
@@
with the explicit form of vertex $\Gamma^{(3)}_{\phi_{q}\phi_{c}\phi_{c},k}$ shown in App. A. The flow equation of $V^{\prime}_{k}(\rho)$ is the same as Model A with one component of the order parameter.
@@
$\partial_{\tau}u^{\prime}(\bar{\rho})=$ $(-2+\eta_{\phi})u^{\prime}(\bar{\rho})+(-2+d+\eta_{\phi})\bar{\rho}u^{(2)}(\bar{\rho})$
-- $\frac{2\nu_{d}}{d}\left(1-\frac{\eta_{\phi}}{d+2}\right)\left(\frac{3u^{(2)}(\bar{\rho})+2\bar{\rho}u^{(3)}(\bar{\rho})}{(1+u^{\prime}(\bar{\rho})+2\bar{\rho}u^{(2)}(\bar{\rho}))^{2}}\right)\,,$ (22)
$\partial_{\tau}\bar{\lambda}=$ $-(z-4+\eta_{\phi})\bar{\lambda}$
@@
As for the static anomalous dimension $\eta_{\phi}$, we utilize the result with the LPA approximation
@@
Here we neglect the loop diagram correction for the coupling $g_{k}$. The flow for $\bar{g}$ only comes from the running of canonical dimension. It has been proved that there is no correction for the coupling to the one-loop order in the perturbative theory [53, 54]. It has been shown that it is the combination of $\bar{\lambda}\bar{\eta}$ that determines the scaling behavior, not the coefficients $\bar{\lambda}$ and $\bar{\eta}$ separately. Therefore we define a new dimensionless coupling constant
@@
with $\nu_{d}=1/[2^{d}\pi^{d/2}\Gamma(d/2)]$ is the geometric factor corresponding to the angular integration. By combining the flow equations of Eq. 23, 24 and 25, one arrives at
@@
$I_{\eta}=$ $\frac{2}{d}\frac{1}{(1+\bar{m}^{2})^{3}}\int\mathrm{d\Omega_{d}}(1-3\cos^{2}\theta+2\cos^{4}\theta)\,.$ (31)
As for the $I_{\lambda}$, the higher order terms in the Eq. 23 have been neglected. The fixed-point equation of the coupling constant $f$ as well as its solution is readily obtained from Eq. 29.
@@
where the exponents $x_{\lambda}$ and $x_{\eta}$ are the anomalous dimensions, they can be computed using the following relation:
@@ IV Numerical results
Figure 4: Left panel: The global solution for $u^{\prime}(\bar{\rho})$ derived from the fixed-point equation Eq. 22. Right panel: The static anomalous dimension $\eta_{\phi}$ as a function of spatial dimension $d$, including a comparison with the $\epsilon$-expansion result presented in Eq. 34.
Figure 5: The dynamic anomalous dimensions $x_{\lambda}$(left panel) and $x_{\eta}$(right panel) across various spatial dimensions $d$. Results from fRG calculation and $\epsilon$-expansion taken from [53] are summarized.
Figure 6: Dynamical critical exponent $z$ defined in Eq. 8 as a function of the spatial dimension $d$. $\epsilon$-expansion for $\eta_{\phi}$ in Eq. 34 and $x_{\lambda}$ in Eq. 9 are also presented as comparison.
In this work, the derivative of global effective potential $u^{\prime}(\bar{\rho})$ is obtained using the large field expansion method proposed in [64]. By directly solving the fixed-point equation starting from the large field asymptotic, we achieve the global effective potential with high numerical accuracy. The fixed-point value of the coupling constant $f^{*}$ is then determined from the global effective potential.
In Fig. 4 we present the global solution of derivatives of the effective potential $u^{\prime}(\bar{\rho})$ as a function of $\bar{\rho}$ for three different spatial dimensions(left panel). We also show the static anomalous dimension $\eta_{\phi}$ spatial dimension $d$ as a function of spatial dimension $d$ which is extracted from Eq. 27(right panel). The function $u^{\prime}(\bar{\rho})$ demonstrates monotonic behavior with respect to $\bar{\rho}$ and possesses a unique zero, indicating the presence of the Wilson-Fisher fixed point. As the spatial dimension $d$ decreases, the position of zero $\bar{\rho}_{0}$ increases, indicating the necessity of a global solution rather than a local Taylor expansion, as discussed in [64]. In the LPA approximation, the static anomalous dimension follows the form given in Eq. 27, and we compare our results with the $\epsilon$-expansion up to the third order[65].
@@
The value of $\eta_{\phi}$ from fRG computation and $\epsilon$-expansion are comparable when $d\geq 3.5$. We also present the exact value $\eta_{\phi}=0.25$ for $d=2$ with a star in the figure.
In Fig. 5, we investigate the dependence of the dynamic anomalous dimensions $x_{\lambda}$, $x_{\eta}$ on the spatial dimension $d$. The $\epsilon$-expansion results from Eq. 9 are also shown for comparison. The anomalous dimension $x_{\lambda}$, associated with the transport coefficient, shows excellent agreement with the perturbative expansion up to $\epsilon^{2}$, with deviations appearing for spatial dimensions $d<2.8$. In contrast, the behavior of $x_{\eta}$ differs. Results from the fRG calculations and $\epsilon$-expansion are comparable only as the spatial dimension approaches 4. Additionally, our result for $x_{\eta}$ exhibits non-monotonic behavior when $d\lesssim 2.5$.
Finally, the dynamic critical exponent $z$ as a function of spatial dimension $d$ is shown in Fig. 6. The anomalous dimensions $\eta_{\phi}$ and $x_{\lambda}$ are obtained from Fig. 4 and Fig. 5 for our real-time fRG computation, respectively. The $\epsilon$-expansion results for $\eta$ are taken from Eq. 34 and for $x_{\lambda}$ from Eq. 9. Since the order parameter is a conserved quantity, the dynamic critical exponent $z=4$ for spatial dimension $d=4$. Both results are highly consistent, with only minor differences emerging as the spatial dimension approaches $2$.
@@ V Summary and outlook
In this work, we investigate the critical dynamics of Model H, which describes a conserved order parameter field coupled with the transverse momentum density, using the real-time fRG approach. The Langevin equation of Model H incorporates kinetic terms and coupling mode interactions. By integrating the noise terms and introducing the Martin-Siggia-Rose response field, i.e., the “q” type fields, which encode fluctuation effects in the Schwinger-Keldysh field theory, we obtain the RG scale-dependent effective action to address the dynamics with the fRG approach.
@@
Our results for the anomalous dimension $x_{\lambda}$ and and the dynamic critical exponent show excellent consistency with the perturbative $\epsilon$-expansion results from previous studies in the range $2\leq d\leq 4$. Only small deviations emerge as the spatial dimension approaches $2$. However, result for anomalous dimension $x_{\eta}$ from fRG computation and $\epsilon$-expansion are comparable only in a narrow range of spatial dimensions close to $4$. Additionally, $x_{\eta}$ exhibits non-monotonic behavior for $d\lesssim 2.5$. From the previous analysis[30], the dynamics of QCD critical end point and Model H belong to the same dynamic universality class. Our results predict a dynamic critical exponent $z=3.0507$ for $d=3$.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China under Grant Nos. 12175030.
@@ Appendix A Propagators and mode-coupling vertices
As mentioned in Sec. III, in the real-time fRG framework, the definition of propagators correspond to two-point function of the effective action which contain the retarded, advanced and Keldysh components. The general definition of the propagator is
@@
Taking the functional derivative with the different types of fields in above equation. In this work, the specific forms of the propagators are
@@
$iG^{K}_{\phi,k}(q)=\frac{4\lambda_{k}\bm{q}^{2}}{q_{0}^{2}+\left[\lambda_{k}\bm{q}^{2}\left(\bm{q}^{2}\Big{(}1+r_{B}\Big{(}\frac{\bm{q}^{2}}{k^{2}}\Big{)}\Big{)}+m_{\phi,k}^{2}\right)\right]^{2}}\,.$ (36)
@@
For the flow equation of effective potential, the vertex $\Gamma^{(3)}_{\phi_{q}\phi_{c}\phi_{c},k}$ emerge in the right hand of Eq. 14 is
@@
To evaluate the flow equation for the two-point correlation function in Fig. 3, the higher order vertices are needed. They can be easily obtained from the functional derivative of $\Gamma_{k}[\Phi]$ w.r.t the corresponding fields $\phi$(or $j$). In this work, the three-point vertices contribute from the coupling modes read
@@
where the explicit momentum dependence from the ”classical” fields are labelled. And the $\bm{e}_{j}$ is the unit vector along the $\bm{j}$ field direction.

### v2

@@ II Langevin equation for Model H
In Eq. 1 and 2 $\lambda_{0}$ and $\eta_{0}$ are the transport coefficient and shear viscosity, respectively. Their dynamical scaling behaviors are described by the exponents $x_{\lambda}$ and $x_{\eta}$ in the critical regime, and these exponents can also be seen as dynamic anomalous dimensions. The Hamiltonian has the usual form except the term $\bm{j}^{2}/2$ included. Thus the static critical properties of $\phi$ are the same as the Ginzburg-Landau model with Ising symmetry. When the system is in thermal equilibrium, the equilibrium distribution function is given by $e^{-\mathscr{H}}$. The coupling constant $g_{0}$ describes the interaction between the two dynamical modes in model H. The Gaussian noises $\theta$ and $\zeta$ drive the system to thermal equilibrium and their correlation functions are shown in Eq. 4.
@@
$D\equiv\frac{\lambda}{\chi_{\phi}}=\frac{R\,k_{B}T}{\eta\xi}\,,$ (7)
in $d=3$ dimensions. Here $\lambda$ and $\eta$ stand for the renormalized transport coefficient and shear viscosity, respectively; $\chi_{\phi}$ denotes the static susceptibility; $T$ is the temperature and $k_{B}$ the Boltzmann constant; $R$ is a constant. Using the scaling relations as follows
$\chi_{\phi}\sim\xi^{2-\eta_{\phi}},\quad\lambda\sim\xi^{x_{\lambda}},\quad\eta\sim\xi^{x_{\eta}}\,,$ (8)
one arrives at the dynamic critical exponent $z$ for the order parameter
@@
where $\eta_{\phi}$ is the static anomalous dimension. Note that the Kawasaki-Stokes relation in Eq. 7 also leaves us with
$x_{\eta}+x_{\lambda}=1-\eta_{\phi}\,,$ (10)
in $d=3$ dimensions. The results of the exponents $x_{\lambda}$ and $x_{\eta}$ in the $\epsilon$-expansion up to the order of $O(\epsilon^{2})$ are obtained as [53]
@@
In the next section, we will compute the aforementioned exponents using the functional renormalization group approach.
@@ III Model H within the real-time fRG approach
$\Gamma_{k}[\Phi]$ $=i\int\mathrm{d}^{d}x\mathrm{d}t\,\,\Bigg{\{}\phi_{q}\left(\frac{\partial}{\partial t}\phi_{c}+\lambda_{k}\bm{\nabla}^{2}\bm{\nabla}^{2}\phi_{c}\right)$
@@
$-g_{k}\bm{\nabla}\phi_{c}(-\bm{\nabla}^{2}\phi_{c})+2\eta_{k}\bm{\nabla}^{2}\bm{j}_{\bm{q},\beta}\Bigg{]}\Bigg{\}}\,.$ (12)
Here, $(\phi_{c},\phi_{q})$ and $(\bm{j}_{\bm{c},\beta},\bm{j}_{\bm{q},\alpha})$ are the order parameter fields and momentum densities with the subscripts ‘c’ and ‘q’ denoting classical and quantum fields, respectively, see [43, 47] for more details. And subscripts $\alpha,\beta$ denote the component of vector $\bm{j}$, i.e., $\alpha,\beta=1,2,...,d$. Different from the analogue in Eq. 1 and 2, the transport and coupling constants $\lambda_{k},\eta_{k}$ and $g_{k}$ in Eq. 12 also receive corrections from interactions, and they are RG-scale dependent. Note that the order parameter belongs to the Ising universality class, thus presents the same static critical properties. The order parameter can be regarded as a linear combination of the chiral condensate and baryon density in QCD [30]. Moreover, one can see that the powers of gradient in front of $\phi_{c}$ are higher than those in Model A or Model C due to the conservation of the order parameter and momentum. The higher-order scatterings among the $\phi$ fields as well as the mass term of order parameter are encoded in the effective potential.
@@
with the explicit expression of vertex $\Gamma^{(3)}_{\phi_{q}\phi_{c}\phi_{c},k}$ shown in App. A. The flow equation of $V^{\prime}_{k}(\rho)$ is the same as Model A with one component of the order parameter.
@@
$\partial_{\tau}u^{\prime}(\bar{\rho})$
$=$ $(-2+\eta_{\phi})u^{\prime}(\bar{\rho})+(-2+d+\eta_{\phi})\bar{\rho}u^{(2)}(\bar{\rho})$
$-\frac{2\nu_{d}}{d}\left(1-\frac{\eta_{\phi}}{d+2}\right)\left(\frac{3u^{(2)}(\bar{\rho})+2\bar{\rho}u^{(3)}(\bar{\rho})}{(1+u^{\prime}(\bar{\rho})+2\bar{\rho}u^{(2)}(\bar{\rho}))^{2}}\right)\,,$ (24)
$\partial_{\tau}\bar{\lambda}$
$=$ $-(z-4+\eta_{\phi})\bar{\lambda}$
@@
As for the static anomalous dimension $\eta_{\phi}$, we utilize the result with the modified local potential approximation (LPA)
@@
Figure 4: Left panel: Global solution for $u^{\prime}(\bar{\rho})$ derived from the fixed-point equation Eq. 24. Right panel: Static anomalous dimension $\eta_{\phi}$ as a function of the spatial dimension $d$, in comparison with the $\epsilon$-expansion result presented in Eq. 36.
Figure 5: Dynamic anomalous dimensions $x_{\lambda}$ (left panel) and $x_{\eta}$ (right panel) as functions of the spatial dimensions $d$. Results from the fRG and $\epsilon$-expansion in [53] are compared. Moreover, the result of $x_{\eta}$ obtained by Ohta and Kawasaki in [64] is also shown in the right panel in blue dashed line, see text for more details.
Here we neglect the loop diagram correction for the coupling $g_{k}$. The flow for $\bar{g}$ only comes from the running of canonical dimension. It has been proved that there is no correction for the coupling to the one-loop order in the perturbative theory [53, 54]. Besides, the relevant symmetries of the system also protect the coupling constant from renormalization. It has been shown that it is the combination of $\bar{\lambda}\bar{\eta}$ that determines the scaling behavior, rather than the coefficients $\bar{\lambda}$ and $\bar{\eta}$ separately. Therefore we define a new dimensionless coupling constant
@@
with $\nu_{d}=1/[2^{d}\pi^{d/2}\Gamma(d/2)]$ is the geometric factor corresponding to the angular integration. By combining the flow equations of 25, 26 and 27, one arrives at
@@
$I_{\eta}=$ $\frac{2}{d}\frac{1}{(1+\bar{m}^{2})^{3}}\int\mathrm{d\Omega_{d}}(1-3\cos^{2}\theta+2\cos^{4}\theta)\,,$ (33)
where the high-order terms in Eq. 32 from Eq. 25 have been neglected. The fixed-point equation of the coupling constant $f$ as well as its solution is readily obtained from Eq. 31.
@@
where the exponents $x_{\lambda}$ and $x_{\eta}$ are the anomalous dimensions, and they can be computed using the following relation:
@@ IV Numerical results
Figure 6: Dynamical critical exponent $z$ defined in Eq. 9 as a function of the spatial dimension $d$. The $\epsilon$-expansion for $\eta_{\phi}$ in Eq. 36 and $x_{\lambda}$ in Eq. 11 are also presented for comparison.
In this work, the derivative of global effective potential $u^{\prime}(\bar{\rho})$ is obtained by using the large field expansion method proposed in [65]. By directly solving the fixed-point equation starting from the large field asymptotic, we achieve the global effective potential with high numerical accuracy. The fixed-point value of the coupling constant $f^{*}$ is then determined from the global effective potential.
In the left panel of Fig. 4 we present the global solution of the first-order derivative of the effective potential $u^{\prime}(\bar{\rho})$ as a function of $\bar{\rho}$ for three different spatial dimensions. The relevant static anomalous dimension $\eta_{\phi}$ in LPA in Eq. 29 is shown in the right panel of Fig. 4. The function $u^{\prime}(\bar{\rho})$ demonstrates monotonic behavior with respect to $\bar{\rho}$ and possesses a unique zero crossing point, indicating the presence of the Wilson-Fisher fixed point. As the spatial dimension $d$ decreases, the position of zero point $\bar{\rho}_{0}$ increases, indicating the necessity of a global solution rather than a local Taylor expansion, as discussed in [65]. In the LPA approximation, the static anomalous dimension follows the form given in Eq. 29, and we compare our results with the $\epsilon$-expansion up to the third order [66].
@@
The value of $\eta_{\phi}$ from fRG computation and $\epsilon$-expansion are comparable when $d\geq 3.5$. We also present the exact value $\eta_{\phi}=0.25$ for $d=2$ with a star in this plot.
In Fig. 5, we investigate the dependence of the dynamic anomalous dimensions $x_{\lambda}$, $x_{\eta}$ on the spatial dimension $d$. The $\epsilon$-expansion results from Eq. 11 are also shown for comparison. The anomalous dimension $x_{\lambda}$, associated with the transport coefficient, shows excellent agreement with the perturbative expansion up to $\epsilon^{2}$, with deviations appearing for spatial dimensions $d<2.8$. In contrast, the behavior of $x_{\eta}$ differs. Results from the fRG calculations and $\epsilon$-expansion are comparable only as the spatial dimension approaches $d=4$. Additionally, our result for $x_{\eta}$ exhibits non-monotonic behavior when $d\lesssim 2.5$. In [64], the exponent for $x_{\eta}$ is evaluated within the formulation of self-consistent mode coupling, where it is found
$x_{\eta}=$ $-(2-\eta_{\phi})^{2}\Gamma(1-\frac{\eta_{\phi}}{2})\Gamma(d-1-\frac{x_{\eta}}{2}+\frac{\eta_{\phi}}{2})$
×$\times\Gamma(2+\frac{x_{\eta}}{2})\Big{\{}4(d-1)\Gamma(\frac{d}{2}+2)\Gamma(2-\frac{d}{2}+\frac{x_{\eta}}{2}-\frac{\eta_{\phi}}{2})$
×$\times\Gamma(\frac{d}{2}-1-\frac{x_{\eta}}{2})\Gamma(\frac{d}{2}+\frac{\eta_{\phi}}{2})\Big{\}}^{-1}\,.$ (37)
The relevant result is shown in the right panel of Fig. 5 in blue dashed line, obtained through numerical iterations for Eq. 37, where the anomalous dimension $\eta_{\phi}$ are used from fRG calculation. Our result is comparable with this result as $d\gtrsim 3$.
Finally, the dynamic critical exponent $z$ as a function of spatial dimension $d$ is shown in Fig. 6. The anomalous dimensions $\eta_{\phi}$ and $x_{\lambda}$ are obtained from Fig. 4 and Fig. 5 for our real-time fRG computation, respectively. The $\epsilon$-expansion results are taken from Eq. 36 for $\eta$ and from Eq. 11 for $x_{\lambda}$. Since the order parameter is a conserved quantity, the dynamic critical exponent $z=4$ for spatial dimension $d=4$. Both results are highly consistent, with only minor differences emerging as the spatial dimension approaches $d=2$.
@@ V Summary and outlook
In this work, we investigate the critical dynamics of Model H, which describes a conserved order parameter field coupled with the transverse momentum density, using the real-time fRG approach. The Langevin equation of Model H incorporates kinetic terms and mode-coupling interactions. By integrating the noise terms and introducing the Martin-Siggia-Rose response field, i.e., the “q” type fields, which encode fluctuation effects in the Schwinger-Keldysh field theory, we obtain the RG scale-dependent effective action within the fRG approach. This formalism is well-suited for investigating the critical dynamics of Model H.
@@
Our results for the anomalous dimension $x_{\lambda}$ and the dynamic critical exponent show excellent agreement with the perturbative $\epsilon$-expansion results in the range $2\leq d\leq 4$. Only small deviations emerge as the spatial dimension approaches $d=2$. However, result for anomalous dimension $x_{\eta}$ from fRG computation and $\epsilon$-expansion are comparable only in a narrow range of spatial dimensions close to $d=4$. Additionally, $x_{\eta}$ exhibits a non-monotonic behavior for $d\lesssim 2.5$. From the analysis in [30], the dynamics of QCD critical end point and Model H belong to the same dynamic universality class. Our results predict a dynamic critical exponent $z=3.0507$ for $d=3$.
After this work is finished, we note that the critical dynamics of Model H is also studied within the real-time fRG independently in [67].
@@ Acknowledgements.
We thank Jan M. Pawlowski, Fabian Rennecke and Vladimir V. Skokov for insightful comments and discussions. This work is supported by the National Natural Science Foundation of China under Grant Nos. 12175030.
@@ Appendix A Propagators and mode-coupling vertices
As mentioned in Sec. III, in the real-time fRG framework, the propagators corresponding to the inverse of the two-point functions of the effective action, contain the retarded, advanced and Keldysh components. The general definition of the propagator is
@@
Taking the functional derivative with respect to the different types of fields, one arrives at the specific forms of the propagators as follows
@@
$iG^{K}_{\phi,k}(q)=\frac{4\lambda_{k}\bm{q}^{2}}{q_{0}^{2}+\left[\lambda_{k}\bm{q}^{2}\left(\bm{q}^{2}\Big{(}1+r_{B}\Big{(}\frac{\bm{q}^{2}}{k^{2}}\Big{)}\Big{)}+m_{\phi,k}^{2}\right)\right]^{2}}\,,$ (39)
@@
For the flow equation of effective potential, the vertex $\Gamma^{(3)}_{\phi_{q}\phi_{c}\phi_{c},k}$ emerging in the right hand of Eq. 16 is
@@
To evaluate the flow equation for the two-point correlation function in Fig. 3, higher-order vertices are needed. They can be easily obtained from the functional derivative of $\Gamma_{k}[\Phi]$ w.r.t the corresponding fields $\phi$(or $j$). In this work, the three-point vertices arising from the coupling modes read
@@
where the explicit momentum dependence from the “classical” fields are labelled. And the $\bm{e}_{j}$ denotes the unit vector along the $\bm{j}$ field direction.